For the data below (1) Determine the degree of relationship, and (2) the percent of variability...

  1. For the data below (1) Determine the degree of relationship, and (2) the percent of variability accounted for. (3) Graph the data and (4) find the line of best fit (find the equation of the line and draw it on the graph).

X

Y

1

2

2

3

4

2

5

6

8

7
Answer #1

Ans:

x y xy x^2 y^2
1 1 2 2 1 4
2 2 3 6 4 9
3 4 2 8 16 4
4 5 6 30 25 36
5 8 7 56 64 49
Total 20 20 102 110 102

a)

correlation coefficient,r=(5*102-20*20)/SQRT((5*110-20^2)*(5*102-20^2))=0.856

There is strong positive relationship between x and y.

b)

percent of variability accounted for,R^2=0.856^2=0.733 or 73.3%

c)

d)

slope,b=(5*102-20*20)/(5*110-20^2)=0.733

intercept,a=(20-0.7333*20)/5=1.0667

Regression equation:

y'=0.7333 x+1.0667